OpenStudy (anonymous):

find the derivative of /[/3 sqrt{x- frac{1/x}}/]/? pls hlp

OpenStudy (anonymous):

\[3*\sqrt{\frac{x-1}{x}}\]???

OpenStudy (anonymous):

To do fractions, type "frac{numerator}{denominator}"

OpenStudy (anonymous):

\[find the derivative of 3\sqrt{x-1/x}\]

OpenStudy (anonymous):

Factor out constants and use the chain rule and you get \[(3 (1+1/x^2))/(2 \sqrt(-1/x+x))\]

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

Natalie, I think you're making it a little more complicated than it needs to be. Just use chain rule: Derivative of\[3\sqrt(x)\]?

OpenStudy (anonymous):

That's the outside function. The inside function is x-(1/x). Chain rule says: do derivative of the ouside function. Multiply by the derivative of the inside function.

OpenStudy (anonymous):

\[3 d/dxsqrt{x}= 3/2sqrt{x}\]\

OpenStudy (anonymous):

\[d/dx \sqrt{1/x}=\]

OpenStudy (anonymous):

ok so the answer s\[3/2\sqrt{x}+1/x^{2}\]

OpenStudy (anonymous):

is this correct?